OpenStudy (anonymous):

How do you find the critical points of f(x)=x+(4/x)

OpenStudy (anonymous):

I found the frst derivative but I don't know what to do after that

OpenStudy (anonymous):

first derivative is equal to zero and solve for x

OpenStudy (anonymous):

once you found the critical x values, plug back in and to get critical points

OpenStudy (anonymous):

I don't know how to break down \[1-\frac{ 4 }{ x^2 }\]

OpenStudy (anonymous):

\[1-\frac{4}{x^2}=0\] solve for x

OpenStudy (anonymous):

oh wait I just got x=-2,2

OpenStudy (anonymous):

ok now substitute them back into the original equation the critical points are?

OpenStudy (anonymous):

-2, 0 2,4

OpenStudy (anonymous):

sorry meant -2,-4

OpenStudy (anonymous):

is that right(-2,-4)(2,4) for the critical points

OpenStudy (anonymous):

Now that I have my critical points can you help me with the second part of my quesiton?